OpenStudy (insa):

a solid metal rocket is made by fastening a cone of vertical height 49cm and base radius 18cm to a circular cylinder of length 192cm and radius 18cm. if the mass of the model is 2145kg, calculate the density of the metal in kg/m^3, giving your answer correct to the nearest whole number.

OpenStudy (anonymous):

Hey. I will be giving you the way of doing this problem. You know the formula of volume of a cone and cylinder.??

OpenStudy (insa):

yup

OpenStudy (anonymous):

formula of volume of a cone and cylinder= V=1/3b.h

OpenStudy (insa):

@shereenkhan i know the formula i have the problem just with calculating the density

OpenStudy (anonymous):

okay.

OpenStudy (paki):

Volume of cone V= 1/3 * pi*r^2*h.... V= 1/3 * pi*(0.18)^2*0.49 V= 0.0166 m^3 ===> here first convert height and base into metre, from cm first, as we need our answer in kg/m3....

OpenStudy (paki):

@insa

OpenStudy (paki):

now find the Volume of cylinder, by the relation... = pi*r^2*h...

OpenStudy (paki):

volume= pi*r^2*h... Volume = pi*(0.18)^2*1.92 V= 0.195 m^3

OpenStudy (anonymous):

i think the density of the metal is 10156.25 kg/m^3.

OpenStudy (insa):

i knw this, i have just the problem with density how to calculate that?

OpenStudy (anonymous):

density = mass/volume. here it is 2145kg/.2112m^3=10156.25kg/m^3

OpenStudy (insa):

thx everyone

OpenStudy (paki):

pleasure... @insa